Bluefield has decided to use an ABC costing system and has identified the following detailed information about its cost pools and cost drivers:

Activity cost pools
Materials handling $59,690
Machine maintenance $17,000
Cost drivers
Number of material moves 470
Number of machine hours 50,000


Required:
1. Calculate Bluefield’s activity rate for each cost pool.
2. Determine the amount of overhead assigned to Bluefield’s products if each has the following activity demands:

Product A Product B
Number of material moves 240 230
Number of machine hours 17,900 32,100

Calculate Bluefield’s activity rate for each cost pool. (Round your final answer for per material moves to nearest dollar amount and per machine hour to 4 decimal places.)

Activity Rate
Materials Handling per Material Move
Machine Maintenance per Machine Hour

Determine the amount of overhead assigned to Bluefield’s products if each has the following activity demands: (Round intermediate calculations to 4 decimal places and your final answers to 2 decimal places.)

Product A Product B
Number of material moves 240 230
Number of machine hours 17,900 32,100

Determine the amount of overhead assigned to Bluefield’s products if each has the following activity demands: (Round intermediate calculations to 4 decimal places and your final answers to 2 decimal places.)

Product A Product B
Number of material moves 240 230
Number of machine hours 17,900 32,100
Total Overhead Assigned
Product A
Product B
Answer #1
1
Activity Rate
Materials Handling 127 per Material Move =59690/470
Machine Maintenance 0.3400 per Machine Hour =17000/50000
2
Total Overhead Assigned
Product A 36566 =(240*127)+(17900*0.34)
Product B 40124 =(230*127)+(32100*0.34)